OpenStudy (anonymous):

solve x^2 -4x +4 = 0

OpenStudy (anonymous):

how would you start?

OpenStudy (anonymous):

-2 2 -2 or 2 0

OpenStudy (anonymous):

i think its the third option

OpenStudy (anonymous):

can you factor it?

OpenStudy (anonymous):

I have already worked it out.. and got (x-2)^2 = 0

OpenStudy (anonymous):

@jeremyggg

OpenStudy (anonymous):

so, what number makes that true? theres your answer!

OpenStudy (anonymous):

Well thats where i need help haha

OpenStudy (anonymous):

is it 2?

OpenStudy (anonymous):

if you tried -2 you'd get (-2-2)^2= (-4)^2 = 16. 16 does not = 0